XYZ Corporation is considering the acquisition of a new piece of equipment. The new equipment costs $250,000, plus it will cost $10,000 to ship it to XYZ, and it will cost $5,000 to install it at XYZ. The equipment has a 10-year class life and will be depreciated by the MACRS depreciation system. What amount of depreciation will be expensed in Year 2?
Answer #1

Depreciable cost = cost of equipment+ shipping cost + installation cost

= 250,000+10,000+5,000 =$265,000

Depreciation using MACRS for year 2 = cost*applicable rate

= 265,000*18%

=$47,700